OpenStudy (anonymous):

Standard form of equation of hyperbola with given characteristics: A) vertices (1,2) and (1,-2) and passes through the point (0,√5)

OpenStudy (anonymous):

The center is directly between the two vertices, meaning that the center is at 1,0

OpenStudy (anonymous):

Because the two vertices are on the same vertical line, the hyperbola is vertical, meaning that the equation is a variation of \[(y-0)^2 - (x-1)^2 = 1\]. Using the vertices once more, we see that the length of the major axis of the hyperbola must have length 4. This means that the radius has length 2, and 2^2 = 4. \[\frac{ y^2 }{ 4 } - \frac{ (x-1)^2 }{ b^2 } = 1\] Plugging in 0,sqrt 5 into the equation, we have that\[\frac{1}{b^2} = \frac{1}{4}\] so b^2 is 4. Our final answer is: \\[\frac{ y^2 }{ 4 } - \frac{ (x-1)^2 }{ 4 } = 1\]

OpenStudy (anonymous):

Thank you for being a life-saver